# NoiseNCA: Noisy Seed Improves Spatio-Temporal Continuity of Neural Cellular Automata

### v2

@@ Abstract
Neural Cellular Automata (NCA) is a class of Cellular Automata where the update rule is parameterized by a neural network that can be trained using gradient descent. In this paper, we focus on NCA models used for texture synthesis, where the update rule is inspired by partial differential equations (PDEs) describing reaction-diffusion systems. To train the NCA model, the spatio-termporal domain is discretized, and Euler integration is used to numerically simulate the PDE. However, whether a trained NCA truly learns the continuous dynamic described by the corresponding PDE or merely overfits the discretization used in training remains an open question. We study NCA models at the limit where space-time discretization approaches continuity. We find that existing NCA models tend to overfit the training discretization, especially in the proximity of the initial condition, also called ”seed”. To address this, we propose a solution that utilizes uniform noise as the initial condition. We demonstrate the effectiveness of our approach in preserving the consistency of NCA dynamics across a wide range of spatio-temporal granularities. Our improved NCA model enables two new test-time interactions by allowing continuous control over the speed of pattern formation and the scale of the synthesized patterns. We demonstrate this new NCA feature in our interactive online demo. Our work reveals that NCA models can learn continuous dynamics and opens new venues for NCA research from a dynamical systems’ perspective.
@@ Introduction
For decades, Cellular Automata (CA) have served as the foundation for studying and creating life-like systems (Von Neumann et al.,, 1966; Gardner,, 1970), demonstrating how simple local interactions can lead to complex and emergent behaviors (Wolfram and Gad-el Hak,, 2003). Neural Cellular Automata (NCA [^1] In this paper, we use NCA to refer to both Neural Cellular Automata and Neural Cellular Automaton.) extend the classical CA framework by employing a neural network to define the update rule. This neural network can be trained using gradient descent, allowing NCA to learn complex dynamics from data. This has led to promising applications in texture synthesis (Niklasson et al., 2021b, ; Pajouheshgar et al., 2023a, ; Pajouheshgar et al., 2023b, ), modeling morphogenesis (Mordvintsev et al.,, 2020), image generation (Palm et al.,, 2022; Otte et al.,, 2021; Tesfaldet et al.,, 2022), and discriminative tasks (Randazzo et al.,, 2020; Sandler et al.,, 2020).
NCA models operate on a discrete grid of cells, where each cell’s state evolves over time according to a local update rule. The NCA update rule is inspired by Partial Differential Equations (PDEs) that describe reaction-diffusion systems (Turing,, 1990) and can be viewed as an Euler integration of the PDE over a discretization of space and time. A key question surrounding NCA is whether it can truly capture underlying continuous dynamics or simply overfits the discrete representations used during training. We investigate this question by studying NCA behavior across different spatio-temporal granularities and at the limit where space-time discretization approaches continuity. In this work, we focus on NCA models trained for texture synthesis.
We find that existing NCA models, depending on the target texture, can fail to generalize to spatio-temporal granularities different from the training discretization (Niklasson et al., 2021b, ; Pajouheshgar et al., 2023b, ). We hypothesize that the overfitting occurs due to NCA’s reliance on stochastic updates to break the initial symmetry between the cells. We propose a simple, yet effective solution to this problem by removing the stochastic updates from the NCA architecture and initializing the cell states with random uniform noise. We call this approach Noise-NCA.
Through qualitative and quantitative evaluations, we show that Noise-NCA significantly improves the consistency of NCA dynamics across different granularities of space and time discretization. Furthermore, our improved NCA model unlocks two new test-time interaction capabilities: continuous control over the speed of pattern formation and the scale of the synthesized textures. We demonstrate these functionalities in an online demo available at https://noisenca.github.io. Our work reveals NCAs’ ability to learn continuous dynamics, opening new avenues for exploring them from a dynamical systems’ perspective.
@@
Inspired by Alan Turing’s seminal work on Reaction-Diffusion (RD) systems (Turing,, 1990), Mordvintsev et al., (2020) propose Neural Cellular Automata (NCA) as a variant of RD systems. Unlike traditional RD systems that require manual design and parameter tuning to create a desired pattern, NCA models can be trained using gradient descent. The authors draw a parallel between NCA and dynamical systems to motivate their design choices. Niklasson et al., 2021b further strengthen this relationship by formalizing the NCA training as numerically solving a Partial Differential Equation (PDE) defined on the cell states. NCA models have been successfully applied in various domains (Pajouheshgar et al., 2023a, ; Pajouheshgar et al., 2023b, ; Mordvintsev and Niklasson,, 2021; Palm et al.,, 2022; Tesfaldet et al.,, 2022; Randazzo et al.,, 2020; Sandler et al.,, 2020). Despite these advances, most studies employing NCA have merely treated the model as a numerical method for solving the underlying dynamics, with barely any attention paid to whether NCA captures continuous dynamics or overfits discrete representations.
@@
Mordvintsev et al., (2021) first investigate whether the spatio-temporal discretization used to train their differentiable RD system can effectively approximate a continuous PDE. However, their model lacks a directional spatial gradient and only involves local isotropic information. This sets them apart from common NCA models (Mordvintsev et al.,, 2020; Niklasson et al., 2021b, ; Pajouheshgar et al., 2023a, ; Pajouheshgar et al., 2023b, ) where cells perceive their neighborhood in a direction-sensitive manner. Moreover, they do not provide a solution for the cases where the model diverges from the expected behavior under the governing PDE. Niklasson et al., 2021a explore the Growing NCA model (Mordvintsev et al.,, 2020) from the perspective of asynchronicity in the update rule and find that a globally synchronous update scheme negatively impacts the temporal continuity of NCA. However, the range of temporal discretizations examined in their work is very limited, leaving the behavior of NCA at the continuous-time limit as an open question. Furthermore, their findings indicate that NCA models, even with asynchronous updates, are prone to overfitting the training discretization. We extend upon previous analyses by pushing NCA to the continuous space-time limit and introduce a simple yet effective solution to improve the spatio-temporal continuity of NCA models.
@@
where $H,W$ are the height and width of the grid, respectively. We define $\mathrm{S}^{\odot}(x,y,t)$ as a set that contains the state of all cells in the neighborhood of a given cell  [^2] Note that the neighborhood can also contain the cell itself.. The most common neighborhood used in NCA models is the 9-point Moore neighborhood.
@@
As in any Cellular Automata, the NCA update rule is local and space/time invariant. Locality implies that, at each update step, a cell can only be influenced by a limited number of cells, namely its neighboring cells. Space-Time invariance means that all cells follow the same update rule, which does not change over time. Pajouheshgar et al., 2023a suggested that the NCA update rule can be decomposed into perception and adaptation stages.
@@
In this stage, each cell gathers information from its neighborhood to form its perception vector $\mathrm{Z}(x,y,t)$ by applying depth-wise convolution on the state of cells in its neighborhood $\mathrm{S}^{\odot}(x,y,t)$:
@@
where $K_{\textup{id}},K_{\textup{x}},K_{\textup{y}},K_{\textup{lap}}$ denote Identity, Sobel-X, Sobel-Y, and Laplacian filters. Note that the perception stage is non-parametric and convolution filters are frozen during training. These filters can be modified in test time to interactively control the behavior of cells (Niklasson et al., 2021b, ; Pajouheshgar et al., 2023a, ; Pajouheshgar et al., 2023b, ).
@@
Training Scheme
We use the same training strategy to train all three NCA variants: Vanilla-NCA, PE-NCA, and our Noise-NCA. To train the model, we start from an initial state $\mathrm{S}(x,y,t=0)$ and iteratively apply the NCA update rule to obtain the cell states $\mathrm{S}(x,y,t)$ at time $t$. The RGB image is then extracted by taking the first 3 channels of $\mathrm{S}(x,y,t)$ and compared to the target texture using a loss function. We use the appearance loss function proposed by Pajouheshgar et al., 2023b which measures the similarity of two texture images using a pre-trained VGG16 network (Simonyan and Zisserman,, 2015). The gradients from the loss function are backpropagated through time and used to optimize the NCA parameters $(\theta)$. The models are trained at resolution of $H,W=(128,128)$. We use the pool technique proposed by Mordvintsev et al., (2020) to improve the long-term stability of the update rule.
@@
In the Perception stage, the spatial derivatives including $\nabla_{\mathrm{x}},\nabla_{\mathrm{y}}$, and $\nabla^{2}$ are estimated using Sobel filters $K_{\textup{x}},K_{\textup{y}}$ and 9-point discrete Laplacian kernel $K_{\textup{lap}}$, respectively. In the Adaptation stage, the cell states are updated using Euler integration with timestep $\Delta\textup{t}$. Note that the Sobel and Laplacian filters used in the Perception stage need to be divided by the cell size $\Delta\textup{x},\Delta\textup{y}$ to get consistent estimations of the spatial derivatives across different resolutions. From the finite difference perspective, a consistent approximation of spatial gradients using Sobel filters can be achieved using:
@@
We analyze and experiment with two NCA baselines. The first baseline is the model proposed by Niklasson et al., 2021b . Figures 2 and 3 accurately depict its architecture, which we refer to as Vanilla-NCA. The second baseline is adapted from the DyNCA model by Pajouheshgar et al., 2023b . We refer to it as PE-NCA where PE is short for positional encoding. PE-NCA appends the normalized cell coordinates $(\frac{x}{H},\frac{y}{W})$ to the perception vector $Z(x,y,t)$, which allows the cells to be aware of their position and thus making $\textup{W}_{1}$ a $(D\times 4C+2)$ matrix. Note that Vanilla-NCA uses circular padding when applying convolution kernels in the perception stage, while PE-NCA uses replicate padding to make the boundary condition consistent with positional encoding. In both of these baselines, the initial state [^4] Also called seed, is set to zero $\mathrm{S}(x,y,t=0)=0.0$. The Vanilla-NCA model relies only on the stochastic update mask $\delta(x,y,t)$ to break the symmetry between cells in the initial state $t=0$, while the positional encoding in PE-NCA acts as another source to differentiate between cells.
@@
We propose a simple yet effective change to the baseline NCAs that will prevent the model from overfitting to the training discretization. Our proposed NCA model, namely Noise-NCA is very similar to the Vanilla-NCA with the main difference being that we use random uniform noise as the initial state, i.e., seed
@@
We train the three NCA variants - Vanilla-NCA, PE-NCA, Noise-NCA - on 45 different textures taken from Pajouheshgar et al., 2023b using the training scheme described previously. We conduct qualitative and quantitative experiments to evaluate the trained NCA models at continuous-time limit $\Delta\textup{t}\to 0^{+}$ and continuous-space limit $\Delta\textup{x},\Delta\textup{y}\to 0$. In the paper, the qualitative results are demonstrated for two textures: ”bubbly_0101”, and ”grid_0135”. The complete qualitative results, which compare different NCA variants on all 45 textures, are available at https://noisenca.github.io/supplementary/. For a quantitative evaluation, we train and test the models on all of the 45 textures and use the appearance loss proposed by Pajouheshgar et al., 2023b as a metric to quantify the quality of the NCA output for different values of $\Delta\textup{t}$ and $\Delta\textup{x},\Delta\textup{y}$.
@@
Given a $\Delta\textup{t}\leq 1.0$ and a simulation time $T$, we start from the initial state $\mathrm{S}(x,y,t=0)$ and use Euler integration to find $\mathrm{S}(x,y,t=T)$ by applying the NCA update rule for $\left\lceil\frac{T}{\Delta\textup{t}}\right\rceil$ steps. We set $T=300$ in our experiments. Figure 4 shows the output of all three NCA variants on two different textures for $\Delta\textup{t}\in\{1.0,0.4,0.1\}$. As shown in the figure, when $\Delta\textup{t}=1.0$ all models successfully create the correct patterns. However, as we decrease $\Delta\textup{t}$ the Vanilla-NCA and PE-NCA fail to synthesize the correct texture. On the contrary, Noise-NCA is able to successfully generate the correct pattern across all values of $\Delta\textup{t}$.
Figure 4: Output of NCA models for different values of $\Delta\textup{t}$ used in the Euler integration on two different textures. As $\Delta\textup{t}$ decreases, the patterns created by Vanilla-NCA and PE-NCA start to deviate from the correct pattern, while Noise-NCA is able to maintain the correct pattern across all values of $\Delta\textup{t}$.
For a more rigorous comparison, we perform a quantitative evaluation on 45 different textures for $\Delta\textup{t}$ values ranging from $[10^{-3},10^{0}]$. Let $\mathrm{S}^{\Delta\textup{t}}_{T}$ denote the output of NCA at time $T$ when $\Delta\textup{t}$ is used as the integration timestep. For each value of $\Delta\textup{t}$, we evaluate
@@
Figure 6 shows the output of all three NCA variants on two different textures for $\Delta\textup{x}\in\{1.0,0.5,0.25\}$. The qualitative results in the figure show that Noise-NCA is able to generate the correct pattern across all values of $\Delta\textup{x}$, while the baseline NCAs fail to generalize to different spatial resolutions. For a quantitative evaluation of the effect of $\Delta\textup{x}$ on the output quality, we use $\Delta\textup{x}$ values ranging from $[2^{-4},2^{0}]$, which results in grids with resolutions varying from $2048\times 2048$ to $128\times 128$ and test the models on 45 different textures. For a given value of $\Delta\textup{x}$, let $\mathrm{S}^{\Delta\textup{x}}_{T}$ denote the output of NCA at time $T$ given spatial discretization $\Delta\textup{x}$ and let $\downarrow_{128}$ denote the bilinear interpolation operator for resizing to resolution $128\times 128$. As a measure of relative output quality, we evaluate the ratio of appearance losses
@@
In this section, we further investigate the NCA model and provide insights into the source of the overfitting. We also study the behavior of NCAs from a fixed point perspective to understand the output of overfitting NCA models at the limit $\Delta\textup{t}\to 0$.
@@
Our experiments in the previous section demonstrate that Vanilla-NCA ovefits to the timestep used during training, namely $\Delta\textup{t}=1.0$. We analyze this phenomenon and provide intuition that motivates and supports the design of Noise-NCA. We introduce two different scheduling policies $\Delta\textup{t}_{\mathrm{A}}(t),\Delta\textup{t}_{\mathrm{B}}(t)$ that will determine the timestep value $\Delta\textup{t}$ as a function of time $t$.
@@
We find that at the continuous-time limit $\Delta\textup{t}\to 0^{+}$, the behavior of the Vanilla-NCA model, either overfit or generalize, largely depends on the target texture. As shown in the second row of Figure 9, when $\Delta\textup{t}$ is small, for textures in Group B, the model is able to create the correct pattern, while for textures in Group A, it converges to a uniform solution that resembles the background color of the target pattern. We speculate that most target textures that have a clear background and foreground separation fall into Group A. In this section, we provide a fixed point analysis of the Vanilla-NCA model to better understand the overfitting problem and the nature of the uniform solution that Vanilla-NCA converges to. In this uniform solution, all cells converge to the same state and become stable. We refer to this state as a fixed point state $\mathrm{S}_{fp}\in\mathbb{R}^{C}$ of the NCA. In the CA literature, this state $\mathrm{S}_{fp}$ is often referred to as a Quiescent State (Sayama,, 2015). The property of the quiescent state is that if all neighbors of a cell, including the cell itself, are in the quiescent state, then the cell will remain in the quiescent state in the next iteration  [^5] For example in the Game of Life CA, dead states are quiescent states since a dead cell remains dead if all of its neighbors are dead.. In the context of NCA, the quiescent state is a local fixed point of the NCA model.
@@
If $\mathrm{S}_{fp}$ is a quiescent state of the NCA model, then the value of $\Delta\mathrm{S}$ in Equation 2 should be close to zero. Let $\mathrm{Z}_{fp}=\left[\mathrm{Z}_{\textup{id}},\mathrm{Z}_{\textup{x}},\mathrm{Z}_{\textup{y}},\mathrm{Z}_{\textup{lap}}\right]$ be the perception vector of a cell in the quiescent state. Since all cells in the neighborhood are in the quiescent state, we can simplify the equation by using the fact that applying the Sobel and Laplacian filters to a constant neighborhood will give zero results. Therefore, $\mathrm{Z}_{\textup{x}}=\mathrm{Z}_{\textup{y}}=\mathrm{Z}_{\textup{lap}}=0$ and $\mathrm{Z}_{\textup{id}}=\mathrm{S}_{fp}$. Let $\left[\textup{W}_{1}^{\textup{id}},\textup{W}_{1}^{\textup{x}},\textup{W}_{1}^{\textup{y}},\textup{W}_{1}^{\textup{lap}}\right]$ be a partitioning of the weight vector $\textup{W}_{1}$ where each partition is a $D\times C$ matrix that operates on its corresponding part of the perception vector. To find a fixed point state of an NCA model we propose to solve the following optimization problem which minimizes the $L_{1}$ norm of the cell update $\Delta\mathrm{S}$:
$\mathrm{S}_{fp}=\operatorname*{arg\,min}_{\mathrm{S}}\left\|\textup{W}_{2}\left(\textup{W}_{1}^{\textup{id}}\mathrm{S}+\textup{b}_{1}\right)_{+}\right\|_{1}$ (8)

### v3

@@ Abstract
Neural Cellular Automata (NCA) is a class of Cellular Automata where the update rule is parameterized by a neural network that can be trained using gradient descent. In this paper, we focus on NCA models used for texture synthesis, where the update rule is inspired by partial differential equations (PDEs) describing reaction-diffusion systems. To train the NCA model, the spatio-temporal domain is discretized, and Euler integration is used to numerically simulate the PDE. However, whether a trained NCA truly learns the continuous dynamic described by the corresponding PDE or merely overfits the discretization used in training remains an open question. We study NCA models at the limit where space-time discretization approaches continuity. We find that existing NCA models tend to overfit the training discretization, especially in the proximity of the initial condition, also called ”seed”. To address this, we propose a solution that utilizes uniform noise as the initial condition. We demonstrate the effectiveness of our approach in preserving the consistency of NCA dynamics across a wide range of spatio-temporal granularities. Our improved NCA model enables two new test-time interactions by allowing continuous control over the speed of pattern formation and the scale of the synthesized patterns. We demonstrate this new NCA feature in our interactive online demo. Our work reveals that NCA models can learn continuous dynamics and opens new venues for NCA research from a dynamical system’s perspective.
@@ Introduction
For decades, Cellular Automata (CA) have served as the foundation for studying and creating life-like systems (Von Neumann et al.,, 1966; Gardner,, 1970), demonstrating how simple local interactions can lead to complex and emergent behaviors (Wolfram and Gad-el Hak,, 2003). Neural Cellular Automata (NCA) extend the classical CA framework by employing a neural network to define the update rule. This neural network can be trained using gradient descent, allowing NCA to learn complex dynamics from data. This has led to promising applications in texture synthesis (Niklasson et al., 2021b, ; Pajouheshgar et al.,, 2024, 2023), modeling morphogenesis (Mordvintsev et al.,, 2020), image generation (Palm et al.,, 2022; Otte et al.,, 2021; Tesfaldet et al.,, 2022), and discriminative tasks (Randazzo et al.,, 2020; Sandler et al.,, 2020).
NCA [^1] In this paper, we use NCA to refer to both Neural Cellular Automata and Neural Cellular Automaton. models operate on a discrete grid of cells, where each cell’s state evolves over time according to a local update rule. The NCA update rule is inspired by Partial Differential Equations (PDEs) that describe reaction-diffusion systems (Turing,, 1990) and can be viewed as an Euler integration of a PDE over a discretization of space and time. A key question surrounding NCA is whether it can truly capture underlying continuous dynamics or simply overfits the discrete representations used during training. We investigate this question by studying NCA behavior across different spatio-temporal granularities and at the limit where space-time discretization approaches continuity. In this work, we focus on NCA models trained for texture synthesis.
We find that existing NCA models (Niklasson et al., 2021b, ; Pajouheshgar et al.,, 2023), depending on the target texture, can fail to generalize to spatio-temporal granularities that are different from the training discretization. We hypothesize that the overfitting occurs due to NCA’s reliance on stochastic updates to break the initial symmetry between the cells. We propose a simple, yet effective solution to this problem by removing the stochastic updates from the NCA architecture and initializing the cell states with random uniform noise. We call this approach Noise-NCA.
Through qualitative and quantitative evaluations, we show that Noise-NCA significantly improves the consistency of NCA dynamics across different granularities of space and time discretization. Furthermore, our improved NCA model unlocks two new test-time interaction capabilities: continuous control over the speed of pattern formation and the scale of the synthesized textures, as shown in Figure 1. We demonstrate these functionalities in an online demo available at https://noisenca.github.io. Our work reveals NCA’s ability to learn continuous dynamics, opening new avenues for exploring them from a dynamical system’s perspective.
@@
Inspired by Alan Turing’s seminal work on Reaction-Diffusion (RD) systems (Turing,, 1990), Mordvintsev et al., (2020) propose Neural Cellular Automata as a variant of RD systems. Unlike traditional RD systems that require manual design and parameter tuning to create a desired pattern, NCAs can be trained using gradient descent. The authors draw a parallel between NCA and dynamical systems to motivate their design choices. Niklasson et al., 2021b further strengthen this relationship by formalizing the NCA training as numerically solving a PDE defined on the cell states. NCA models have been successfully applied in various domains (Pajouheshgar et al.,, 2024, 2023; Mordvintsev and Niklasson,, 2021; Palm et al.,, 2022; Tesfaldet et al.,, 2022; Randazzo et al.,, 2020; Sandler et al.,, 2020). Despite these advances, most studies employing NCA have barely paid any attention to whether the model truly captures continuous dynamics or overfits to the training discrete representation.
@@
Mordvintsev et al., (2021) first investigate whether the spatio-temporal discretization used to train their differentiable RD system can effectively approximate a continuous PDE. However, their model lacks a directional spatial gradient and only involves local isotropic information. This sets them apart from common NCA models (Mordvintsev et al.,, 2020; Niklasson et al., 2021b, ; Pajouheshgar et al.,, 2024, 2023) where cells perceive their neighborhood in a direction-sensitive manner. Moreover, they do not provide a solution for the cases where the model diverges from the expected behavior under the governing PDE. Niklasson et al., 2021a explore the Growing NCA model (Mordvintsev et al.,, 2020) from the perspective of asynchronicity in the update rule and find that a globally synchronous update scheme negatively impacts the temporal continuity of NCA. However, the range of temporal discretizations examined in their work is very limited, leaving the behavior of NCA at the continuous-time limit as an open question. Furthermore, their findings indicate that NCA models, even with asynchronous updates, are prone to overfitting the training discretization. We extend upon previous analyses by pushing NCA to the continuous space-time limit and introduce a simple yet effective solution to improve the spatio-temporal continuity of NCA models.
@@
where $H,W$ are the height and width of the grid, respectively. We define $\mathrm{S}^{\odot}(x,y,t)$ as a set that contains the state of all cells in the neighborhood of a given cell  [^2] The neighborhood can also contain the cell itself.. The most common neighborhood used in NCA models is the 9-point Moore neighborhood.
@@
As in any Cellular Automata, the NCA update rule is local and space/time invariant. Locality implies that, at each update step, a cell can only be influenced by a limited number of cells, namely its neighboring cells. Space-Time invariance means that all cells follow the same update rule, which does not change over time. Pajouheshgar et al., (2024) suggested that the NCA update rule can be decomposed into perception and adaptation stages.
@@
In this stage, as shown in Figure 2, each cell gathers information from its neighborhood to form its perception vector $\mathrm{Z}(x,y,t)$ by applying depth-wise convolution on the state of cells in its neighborhood $\mathrm{S}^{\odot}(x,y,t)$:
@@
where $K_{\textup{id}},K_{\textup{x}},K_{\textup{y}},K_{\textup{lap}}$ denote Identity, Sobel-X, Sobel-Y, and Laplacian filters. Note that the perception stage is non-parametric, meaning that the convolution filters are frozen during training. These filters can be modified in test time to interactively control the behavior of cells (Niklasson et al., 2021b, ; Pajouheshgar et al.,, 2024, 2023).
@@
In the Perception stage, the spatial derivatives including $\nabla_{\mathrm{x}},\nabla_{\mathrm{y}}$, and $\nabla^{2}$ are estimated using Sobel filters $K_{\textup{x}},K_{\textup{y}}$ and 9-point discrete Laplacian kernel $K_{\textup{lap}}$, respectively. In the Adaptation stage, the cell states are updated by applying Euler integration with timestep $\Delta\textup{t}$. Note that the Sobel and Laplacian filters used in the Perception stage need to be divided by the cell size $\Delta\textup{x},\Delta\textup{y}$ to get consistent estimations of the spatial derivatives across different resolutions. According to the finite difference method, a consistent approximation of spatial gradients using Sobel filters can be achieved using:
@@
We analyze and experiment with two NCA baselines. The first baseline is the model proposed by Niklasson et al., 2021b . Figures 2 and 3 accurately depict its architecture, which we refer to as Vanilla-NCA. The second baseline is adapted from the DyNCA model by Pajouheshgar et al., (2023). We refer to it as PE-NCA where PE is short for positional-encoding. PE-NCA appends the normalized cell coordinates $(\frac{x}{H},\frac{y}{W})$ to the perception vector $Z(x,y,t)$, which allows the cells to be aware of their position and thus making $\textup{W}_{1}$ a $(D\times 4C+2)$ matrix. Note that Vanilla-NCA uses circular padding when applying convolution kernels in the perception stage, while PE-NCA uses replicate padding to make the boundary condition consistent with positional encoding. In both of these baselines, the initial state [^4] Also called seed, is set to zero $\mathrm{S}(x,y,t=0)=0.0$. The Vanilla-NCA model relies only on the stochastic update mask $\delta(x,y,t)$ to break the symmetry between cells in the initial state $t=0$, while the positional encoding in PE-NCA acts as another source to differentiate between cells.
@@
We propose a simple yet effective change to the baseline NCAs that will prevent the model from overfitting the training discretization. Our proposed NCA model, namely Noise-NCA is similar to the Vanilla-NCA with the main difference being that we use random uniform noise as the initial state, i.e., seed
@@
Training Scheme
We use the same training strategy to train all three NCA variants: Vanilla-NCA, PE-NCA, and our Noise-NCA. To train the model, we start from an initial state $\mathrm{S}(x,y,t=0)$ and iteratively apply the NCA update rule to obtain cell states $\mathrm{S}(x,y,t)$ at time $t$. The RGB image is then extracted by taking the first 3 channels of $\mathrm{S}(x,y,t)$ and compared to the target texture using a loss function. We use the appearance loss function proposed by Pajouheshgar et al., (2023) which measures the similarity of two texture images using a pre-trained VGG16 network (Simonyan and Zisserman,, 2015). The gradients from the loss function are backpropagated through time and used to optimize the NCA parameters $(\theta)$. All models are trained at resolution of $H,W=(128,128)$. We use the pool technique proposed by Mordvintsev et al., (2020) to improve the long-term stability of the update rule.
@@
We train the three NCA variants - Vanilla-NCA, PE-NCA, Noise-NCA - on 45 different textures taken from Pajouheshgar et al., (2023) using the training scheme described previously. We conduct qualitative and quantitative experiments to evaluate the trained NCA models at continuous-time limit $\Delta\textup{t}\to 0^{+}$ and continuous-space limit $\Delta\textup{x},\Delta\textup{y}\to 0$. In the paper, the qualitative results are demonstrated for two textures: ”bubbly_0101”, and ”grid_0135”. The complete qualitative results, which compare different NCA variants on all 45 textures, are available at https://noisenca.github.io/supplementary/. For a quantitative evaluation, we test the models on all of the 45 textures and use the appearance loss proposed by Pajouheshgar et al., (2023) as a metric to quantify the quality of the NCA output for different values of $\Delta\textup{t}$ and $\Delta\textup{x},\Delta\textup{y}$.
@@
Given a $\Delta\textup{t}\leq 1.0$ and a simulation time $T$, we start from the initial state $\mathrm{S}(x,y,t=0)$ and use Euler integration to find $\mathrm{S}(x,y,t=T)$ by applying the NCA update rule for $\left\lceil\frac{T}{\Delta\textup{t}}\right\rceil$ steps. We set $T=300$ in our experiments. Figure 4 shows the output of the three NCA variants on two different textures for $\Delta\textup{t}\in\{1.0,0.4,0.1\}$. As shown in the figure, when $\Delta\textup{t}=1.0$ all models successfully create the correct patterns. However, as we decrease $\Delta\textup{t}$ the Vanilla-NCA and PE-NCA fail to synthesize the correct texture. On the contrary, Noise-NCA is able to successfully generate the correct pattern across all values of $\Delta\textup{t}$.
Figure 4: Output of NCA models for different values of $\Delta\textup{t}$ used in the Euler integration, on two different textures. As $\Delta\textup{t}$ decreases, the patterns created by Vanilla-NCA and PE-NCA start to deviate from the correct pattern, while Noise-NCA is able to maintain the correct pattern for all $\Delta\textup{t}$ values.
For a more rigorous comparison, we perform a quantitative evaluation on 45 different textures for ten different $\Delta\textup{t}$ values ranging from $[10^{-3},10^{0}]$. Let $\mathrm{S}^{\Delta\textup{t}}_{T}$ denote the output of NCA at time $T$ given $\Delta\textup{t}$ as the integration timestep. For each value of $\Delta\textup{t}$, we evaluate
@@
Figure 6 shows the output of all three NCA variants on two different textures for $\Delta\textup{x}\in\{1.0,0.5,0.25\}$. The qualitative results in the figure show that Noise-NCA is able to generate the correct pattern across all values of $\Delta\textup{x}$, while the baseline NCAs fail to generalize to different spatial resolutions. For a quantitative evaluation of the effect of $\Delta\textup{x}$ on the output quality, we use eleven different $\Delta\textup{x}$ values ranging from $[2^{-4},2^{0}]$, which results in grids with resolutions varying from $2048\times 2048$ to $128\times 128$ and test the models on all 45 textures. For a given value of $\Delta\textup{x}$, let $\mathrm{S}^{\Delta\textup{x}}_{T}$ denote the output of NCA at time $T$ given cell size $\Delta\textup{x}$ and let $\downarrow_{128}$ denote the bilinear interpolation operator for resizing to resolution $128\times 128$. As a measure of relative output quality, we evaluate the ratio of appearance losses
@@
In this section, we further investigate NCA and provide insights into the source of the overfitting. We also study the behavior of NCAs from a fixed point perspective to understand the output of overfitting NCA models at the limit $\Delta\textup{t}\to 0$.
@@
Our experiments in the previous section demonstrate that Vanilla-NCA overfits the timestep used during training, namely $\Delta\textup{t}=1.0$. We analyze this phenomenon and provide intuition that motivates and supports the design of Noise-NCA. We introduce two different scheduling policies $\Delta\textup{t}_{\mathrm{A}}(t),\Delta\textup{t}_{\mathrm{B}}(t)$ that will determine the timestep value $\Delta\textup{t}$ as a function of time $t$.
@@
We find that at the continuous-time limit $\Delta\textup{t}\to 0^{+}$, the behavior of the Vanilla-NCA model, either overfitting or generalizing, largely depends on the target texture. As shown in the second row of Figure 9, when $\Delta\textup{t}$ is small, for textures in Group B, the model is able to create the correct pattern, while for textures in Group A, it converges to a uniform solution that resembles the background color of the target pattern. We speculate that this phenomenon is related to the presence of solitons in the NCA output as the textures in Group A exhibit soliton like patterns and have a background/foreground separation. In this section, we provide a fixed point analysis of the Vanilla-NCA model to better understand the overfitting problem and the nature of the uniform solution that Vanilla-NCA converges to. In this uniform solution, all cells converge to the same state and become stable. We refer to this state as a fixed point state $\mathrm{S}_{fp}\in\mathbb{R}^{C}$ of the NCA. In the CA literature, this state $\mathrm{S}_{fp}$ is often referred to as a Quiescent State (Sayama,, 2015). The property of the quiescent state is that if all neighbors of a cell, including the cell itself, are in the quiescent state, then the cell will remain in the quiescent state in the next iteration  [^5] For example in the Game of Life CA, dead states are quiescent states since a dead cell remains dead if all of its neighbors are dead.. In the context of NCA, the quiescent state is a local fixed point of the NCA model.
@@
If $\mathrm{S}_{fp}$ is a quiescent state of the NCA model, then the value of $\Delta\mathrm{S}$ in Equation 2 should be zero. Let $\mathrm{Z}_{fp}=\left[\mathrm{Z}_{\textup{id}},\mathrm{Z}_{\textup{x}},\mathrm{Z}_{\textup{y}},\mathrm{Z}_{\textup{lap}}\right]$ be the perception vector of a cell in the quiescent state. Since all neighboring cells are also in the quiescent state, we can simplify the equation by using the fact that applying the Sobel and Laplacian filters to this neighborhood will yield zero output. Therefore, $\mathrm{Z}_{\textup{x}}=\mathrm{Z}_{\textup{y}}=\mathrm{Z}_{\textup{lap}}=0$ and $\mathrm{Z}_{\textup{id}}=\mathrm{S}_{fp}$. Let $\left[\textup{W}_{1}^{\textup{id}},\textup{W}_{1}^{\textup{x}},\textup{W}_{1}^{\textup{y}},\textup{W}_{1}^{\textup{lap}}\right]$ be a partitioning of the weight vector $\textup{W}_{1}$ where each partition is a $D\times C$ matrix that operates on its corresponding part of the perception vector. To find a fixed point state of an NCA model we propose to solve the following optimization problem which minimizes the $L_{1}$ norm of the cell update vector $\Delta\mathrm{S}$:
$\mathrm{S}_{fp}=\operatorname*{arg\,min}_{\mathrm{S}\in\mathbb{R}^{C}}\left\|\textup{W}_{2}\left(\textup{W}_{1}^{\textup{id}}\mathrm{S}+\textup{b}_{1}\right)_{+}\right\|_{1}$ (8)